OpenStudy (anonymous):

find an equation for the verticle line and the horizontal line through the given point (1, 2/9)

OpenStudy (anonymous):

please explain! (:

jimthompson5910 (jim_thompson5910):

all vertical lines are of the form x = h, where h is any number

jimthompson5910 (jim_thompson5910):

and all horizontal lines are of the form y = k, the k can be any number

jimthompson5910 (jim_thompson5910):

so if you want a vertical line going through (1,2/9), then that vertical line must be x = 1

jimthompson5910 (jim_thompson5910):

since all points on x = 1 have an x coordinate of 1

jimthompson5910 (jim_thompson5910):

I'm sure you can see what the horizontal line must be

OpenStudy (anonymous):

y= 2/9? :)

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

thank you ! :)

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

mind just helping me on a few more? :3

jimthompson5910 (jim_thompson5910):

sure I can help

OpenStudy (anonymous):

how would you write an equation for a line that passes (3,2) and has a slope of -2/3

jimthompson5910 (jim_thompson5910):

what is the general equation for slope intercept form

OpenStudy (anonymous):

y= mx+b?

jimthompson5910 (jim_thompson5910):

what is m? what is b?

OpenStudy (anonymous):

m is the slope and b is the y?

jimthompson5910 (jim_thompson5910):

b is the y _____ fill in the blank

OpenStudy (anonymous):

intercept

jimthompson5910 (jim_thompson5910):

so any line is uniquely determined by its slope and y-intercept

jimthompson5910 (jim_thompson5910):

if you know those two things, you can find the equation of the line

jimthompson5910 (jim_thompson5910):

we have the slope, which is -2/3, so that means m = -2/3

OpenStudy (anonymous):

so y= -2/3x+ 2

jimthompson5910 (jim_thompson5910):

but we don't know the y-intercept

OpenStudy (anonymous):

oh..

jimthompson5910 (jim_thompson5910):

that's if we had (0,2) as a point on the line

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

we have (3,2) instead

jimthompson5910 (jim_thompson5910):

so what we do is plug in x = 3 and y = 2 since that's what (3,2) means

jimthompson5910 (jim_thompson5910):

y = mx + b y = (-2/3)x + b ... plug in the given slope 2 = (-2/3)*(3) + b ... plug in the given point

jimthompson5910 (jim_thompson5910):

notice how every variable but b has been replaced you can now solve for b

jimthompson5910 (jim_thompson5910):

I'll let you do that

OpenStudy (anonymous):

so b=4?

jimthompson5910 (jim_thompson5910):

good, so m = -2/3 (given) and b = 4 (just found out)

jimthompson5910 (jim_thompson5910):

therefore, the equation is \[\large y = -\frac{2}{3}x + 4\]

OpenStudy (anonymous):

:) yay! ok

jimthompson5910 (jim_thompson5910):

glad it's all clicking

OpenStudy (anonymous):

so for the problem what's the equation if the line asses through (-4,9) and has no slope would b=9??

jimthompson5910 (jim_thompson5910):

which type of lines have no slope?

OpenStudy (anonymous):

.__.

OpenStudy (anonymous):

horizontal :D

jimthompson5910 (jim_thompson5910):

no horizontal lines have a slope of 0

OpenStudy (anonymous):

or verticle? ._.

jimthompson5910 (jim_thompson5910):

vertical lines have no slope or an undefined slope

OpenStudy (anonymous):

ok, :) so it's just be y=-4?

jimthompson5910 (jim_thompson5910):

y = -4 is a horizontal line why? because EVERY point on this line has a y coordinate of -4

jimthompson5910 (jim_thompson5910):

so that's not the answer

OpenStudy (anonymous):

i'm confusedd.. :(

jimthompson5910 (jim_thompson5910):

think of two points that have a y coordinate of -4

OpenStudy (anonymous):

(1,-4) (2,-4) ?

jimthompson5910 (jim_thompson5910):

plot those two points and draw a straight line through them

OpenStudy (anonymous):

lol........ >.<

OpenStudy (anonymous):

wow.

jimthompson5910 (jim_thompson5910):

what kind of line did you plot

OpenStudy (anonymous):

horizontal :p

OpenStudy (anonymous):

>.< silly me

jimthompson5910 (jim_thompson5910):

its ok, it's all part of the learning process

jimthompson5910 (jim_thompson5910):

so what you want is a vertical line through (-4, 9)

jimthompson5910 (jim_thompson5910):

since all vertical lines have no slopes

OpenStudy (anonymous):

x=-4....... ?? :3

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

you're a pro already

OpenStudy (anonymous):

Lol NO :(

jimthompson5910 (jim_thompson5910):

yes you are, just keep practicing and you'll get better

OpenStudy (anonymous):

what happens when they don't tell you if it has a slope or not?

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (anonymous):

the question is Write the equation that has a y-intercept -3 and x-intercept -8?

OpenStudy (anonymous):

is that kinda the same as the last one?

jimthompson5910 (jim_thompson5910):

when they say y-intercept -3, what dos that mean?

jimthompson5910 (jim_thompson5910):

does*

OpenStudy (anonymous):

that (_,-3)...... .__. ? lol

jimthompson5910 (jim_thompson5910):

what goes in that blank

OpenStudy (anonymous):

b= -3?

jimthompson5910 (jim_thompson5910):

yes that is true

jimthompson5910 (jim_thompson5910):

but what do you know about a point on this line if the y-intercept is -3

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

a y-intercept of -3 means that the line crosses the y-axis at -3 on the y-axis